OpenStudy (anonymous):

Without using a calculator, find the exact value of cos^-1(cos(17pi/5)). Justify your answer

ganeshie8 (ganeshie8):

without using calculator can u tell the exact value of below expression ? :- \(\large 9876 \times \frac{1}{9876} \pi\)

OpenStudy (anonymous):

isn't that just pi because both 9876 cancel each other ?

OpenStudy (ybarrap):

$$\large \text{Note:} \cos^{-1}(\cos x)= x$$

ganeshie8 (ganeshie8):

true, they cance out cuz division is inverse of multiplication similarly cos^-1 is inverse of cos, so.. ?

ganeshie8 (ganeshie8):

*cancel

OpenStudy (ybarrap):

`"x" here will be exact `

OpenStudy (anonymous):

so they will cancel too and it will become x ??

OpenStudy (anonymous):

or will the answer just be 17pi/5 ?

ganeshie8 (ganeshie8):

\(cos^-1(cos(17\pi/5)) \) \(\cancel{cos^-1}\cancel{cos}(17\pi/5) \) \(17\pi/5\)

ganeshie8 (ganeshie8):

yes !

OpenStudy (anonymous):

thanks ! :)

ganeshie8 (ganeshie8):

np, keep in mind, u need to be bit careful if x is negative.